OpenStudy (hddsjm1za):

The value of acceleration due to gravity (g) on a point 10,000 kilometers above sea level is about 1.49 meters/second2. How much will an object, which weighs 98 newtons on the surface of Earth, weigh on this point? The value of acceleration due to gravity on Earth is 9.8 meters/second2.

OpenStudy (sharmagk83):

mass of the object can be calculated as \[W = mg\] \[m = \frac{ W }{ g }= \frac{ 98 N }{ 9.8 m/s^{2}} = 10 kg\] So weight at a height 10,000 km will be \[W' =mg' = 10 kg \times1.49m/s^{2}= 14.9 N\]